CURRICULUM VITAE

1 PERSONAL DETAILS

Full name  Tomasz Aleksander Kowalewski
Date of Birth  25th September, 1947
Place of Birth  Łódź, Poland
Citizenship  Polish
Married with Teresa Rycerz  1969
Children  Ingrid 1970 - 2004, son Oskar 1973

Work Address Institute of Fundamental Technological Research Polish Academy of Science (IPPT PAN) Pawinskiego 5B, PL 02-106 Warszawa, Poland Phone: +48-228269803, Fax: +48-228269815 e-mail: tkowale@ippt.pan.pl http://www.ippt.pan.pl/en/staff/tkowale

Home Address Sikorskiego 11 m 81, PL 02-758 Warszawa Phone: +48-224074330, +48-602193472 (mobile)

2 ACADEMIC QUALIFICATIONS

Degrees M.Sc. in Physics with Dean’s Distinction Award, Warsaw University, 1969. Ph.D., Institute of Fundamental Technological Research, Polish Academy of Science (IPPT PAN), Warsaw, 1982. Habilitation, Institute of Fundamental Technological Research, Polish Academy of Science (IPPT PAN), Warsaw, 1995. Title of Professor, President of Poland, 2007.

3 APPOINTMENTS

2017-2018 Professor, Deputy Head of the Department of Biosystems and Soft Matter, Institute of Fundamental Technological Research, Polish Academy of Science, Warsaw, Poland.

2010-2017 Professor, Deputy Head of the Department of Mechanics and Physics of Fluids, Institute of Fundamental Technological Research, Polish Academy of Science, Warsaw, Poland.

2009-2013 Deputy Director for Research, Institute of Fundamental Technological Research, Polish Academy of Science, Warsaw, Poland.

2001 – 2009 Head of the Department of Mechanics and Physics of Fluids, Institute of Fundamental Technological Research, Polish Academy of Science, Warsaw, Poland.

1995 – 2001 Head of the Division of Thermal and Free Surface Flows, Center of Mechanics, Institute of Fundamental Technological Research, Polish Academy of Science (IPPT PAN), Warsaw, Poland.

June 1993 – June 1994 Wissenschaftlicher Mitarbeiter (Senior Research Associate), Max-Planck-Institut für Strömungsforschung, Göttingen, Germany

March 1988 – May 1993 Wissenschaftlicher Mitarbeiter (Senior Research Associate), Institut für Angewandte Mechanik und Strömungsphysik, Georg-August-Universität, Göttingen.

Oct. 1985 – Feb. 1988 Visiting Scholar, research fellowship by Max-Planck-Institut, Göttingen.

Sept. 1969 – June 1987 Asystent (Research Assistant) and Adiunkt (Senior Research Associate), Institute of Fundamental Technological Research, Polish Academy of Science (IPPT PAN), Department of Fluid Mechanics, Warsaw, Poland.

4 SCHOLARSHIPS

Aug. – Sept. 1978 Visiting scholar at the Delft University, Holland (by Prof. J.A. Steketee). Hot-Wire Anemometry.

Nov. – Dec. 1980 Visiting scholar at the Max-Planck-Institut für Strömungsforschung, Göttingen (by Prof. E.-A. Müller). Investigations of twophase flow with help of Laser-Doppler-Anemometry.

Mar. – June 1982 Visiting scholar at the Stanford University (by Prof. A. Acrivos). M. Curie-Sklodowska joined Polish-American project on sedimentation and flow of suspensions.

Oct. 1985 – Feb. 1988 Scientific fellowship from the Max-Planck-Society, Institut für Strömungsforschung, Göttingen (by Prof. E.-A. Müller). Experimental modeling of fuel injection and evaporation; fuel film deposition on the walls of engine manifolds.

Feb. – Apr. 1993 Visiting Fellow at the University of New South Wales, School of Mechanical and Manufacturing Engineering. (by Prof. G. de Vahl Davis). Experimental and numerical investigation of the natural convection.

Nov. 1993 Short visit at the Israel Institute of Technology, TECHNION, Dept. of Mechanical Engineering (by Prof. A. Yarin), Haifa, Israel. Investigations of liquid jet instability.

Mar. 1995 – Jan. 1996 Visiting Fellow at the Israel Institute of Technology, Lady Davis Fellowship, TECHNION, Dept. of Aerospace Engineering (by Prof. M. Wolfshtein), Haifa, Israel. Investigations of droplet combustion. Teaching (two semesters course) Advanced Methods in Aerodynamics.

April 1997 Visiting Professor at the University Paris XI (Orsay), lecturing: “Advanced experimental methods in fluid mechanics”.

April-May 1998 Visiting Professor at the University “La Sapienza” and Roma Tre (Rome), application of the PIV method for combustion flow

Nov. 1998 Short visit at the Israel Institute of Technology, TECHNION, Dept. of Mechanical Engineering (by Prof. G. Hetsroni), Haifa, Israel. Joint project on turbulent flow in the channel

March 1999  Visiting Professor at the University Paris XI (Orsay)

7 months within 1999 - 2002 Visiting Scholar at the Arizona State University, Joint atmospheric research, VTMX2000 program

June 2000 Short visit at the Israel Institute of Technology, TECHNION, Dept. of Mechanical Engineering (by Prof. G. Hetsroni), Haifa, Israel. Joint project on turbulent flow in the channel

October 2001 Short visit at the Israel Institute of Technology, TECHNION, Dept. of Mechanical Engineering (by Prof. G. Hetsroni), Haifa, Israel. Joint project on turbulent flow in the channel

March 2002  Visiting Professor at the University Paris XI (Orsay)

Nov. 2002 Short visit at the Israel Institute of Technology, TECHNION, Dept. of Aerospace Engineering (by Prof. Y. Levy), Haifa, Israel. Joint 5FP EU project on combustion FLOXCOM

Mar - April 2003 Visiting Professor at the University of New South Wales (Sydney), CFD Group of Prof. E. Leonardi

Dec 2007 - Jan 2008 Visiting Professor at the University of New South Wales (Sydney), CFD Group of Prof. E. Leonardi

5 RESEARCH EXPERIENCE

1969 – 1985 Research work at the Institute of Fundamental Technological Research (Warsaw), during the experimental study towards Ph.D. degree on low Reynolds number flow of suspensions; experimental investigations of the jet break-up, modeling of the fibre spinning.

1986 – 1987 Research work at the Max-Planck-Institut, Evaporation of fuels and retrograde fluids; atomisation of fuels; investigation of natural convection.

1988 – 1991 Research grants from Deutsche Forschungsgemeinschaft through the University Göttingen: Experimental investigations of the instabilities of the evaporating liquid jets (DFG 201/18) and

1991 – 1993  Dynamic properties of the jets of liquid mixtures (DFG201/20).

1995 – 1996 9 months Lady Davis Fellowship, Application of PIV and image processing for diagnostics of flow with combustion, Israel Institute of Technology, TECHNION, Haifa

1996 – 2010 Research collaboration (project conjoins) IPPT PAN – LIMSI CNRS, L’etude numerique et experimentele de la dynamique du gossissement et du decrochement des bulles de vapeur dans le cas de l’ebullition heterogene

1998 – 2002 Research collaboration IPPT PAN - TECHNION in framework of Polish Academy of Sciences – Israeli Academy of Sciences agreement, on Turbulent thermal heat transfer in a channel flow

1999 – 2002 Research collaboration with Arizona State Univerity (USA) on Enviromental Fluid Dynamics

Other activities Nonlinear dynamic of the oscillating droplet; evaporation of droplets; evaporation induced surface instabilities; image processing; high-Speed video imaging; flow visualisation; tracking of tracer particles in the flow; natural convection; convection with phase changes; Particle Image Velocimetry (PIV), liquid crystals thermography (PIT), computational fluid dynamics, micro-fluidics, microPIV, nanofibres, drug release, Brownian dynamics of nano-objects.

6 MAIN RESEARCH TOPICS

Keywords: Fluid mechanics, micro-fluidics, nanofibres, bioengineering, multiphase flow, free surface flow, thermally driven flow, flow with phase changes. Experimental and numerical methods in fluid mechanics.

• Experimental investigations of liquid jet break-up. Evidence of “super” thin liquid bridges between liquid jet and the droplet. Instability of evaporating jets, modeling of droplet and liquid jet evaporation. Evaporation of emulsions [9, 10, 13, 31, 35, 43-45], microflows [116,119,120], nanofibres [103,113,124,132,133,136,137,140,152,153,155,156].

• Experimental and numerical investigation of nonlinear droplet oscillations. Description of droplet deformation. Development of non-intrusive method to measure dynamic surface tension, measurements of the physical parameters of evaporating surface by oscillating droplet method [11, 18, 20, 21, 25, 26, 36, 37, 40, 42].

• Development of high speed video recording methods based on CCD “frame-transfer” camera and image processor. Application of the method for investigation of droplets and jet instabilities [8, 11, 15, 19, 32].

• Experimental and numerical investigation of natural convection, identification of new flow structures; investigation of natural convection in presence of phase change (freezing) [12, 14, 16, 24, 27, 29, 41, 46-54,67].

• Development of new experimental technique for instantaneous measurement of temperature and velocity two-diemnsional fields in thermal flow, based on application of liquid crystal as tracers [7, 14, 29, 34,48,49,89,94,101,102,121,154].

• Investigations of low Reynolds number two-phase flow in tubes (solid and liquid suspensions). Application of new measuring techniques for measuring velocity and particle concentration, laser doppler and ultrasound doppler anemometry [1-6].

7 EDUCATIONAL EXPERIENCE

Direct supervision of 7 M.Sc. and 3 Ph.D. students graduated at the Georg-August University Göttingen 1985-1994; 3 Ph.D. and 5 M.Sc. students at Warsaw University of Technology. Supervisor of 6 Ph.D. students at IPPT PAN.

Graduate courses:

• Advanced Methods in Aerodynamics (two semesters) given at TECHNION, Dept. of Aerospace Engineering, 1995-1996.

• Advanced Experimental Methods in Fluid Mechanics (professor invite) short courses given at University Paris XI (Orsay), 1997, 1999, 2001.

• Experiment and Modelling Solidification Processes (professor invite) short courses given at University Roma 3 and La Sapienza, 2001.

• Experimental Methods for Quantitative Analysis of Thermally Driven Flows, lecturer and coordinator of the CISM course on Phase Change with Convection, Udine (Italy), 2002.

• Contemporary Experimental Methods in Fluid Mechanics advanced course given at Technical University of Czestochowa, 2011

• Nanofluids and Nanofibres advanced course given at Technical University of Krakow, 2014

8 REVIEWER

Journals Journal of Fluid Mechanics Physics of Fluids Archives of Mechanics (member of Editorial Board) Computer Assisted Mechanics and Engineering Sciences Journal of Engineering Mathematics (member of Board of Assoc. Referees) Journal of Heat Mass Transfer Applied Thermal Engineering JMEMS, Meccanica, APS PRL, APS PRE, AIChEJ Machine Graphics Vision Fluid Dynamics & Materials Processing (member of Editorial Board) Bulletin of Polish Academy of Sciences (member of Editorial Board) Int. J. Thermal Scs. Industrial & Engineering Chemistry Research Exp. in Fluids J. Micromechanics and Microengineering Measurements Science and Technology

Reviewer of Research grants, professor nomination National Science Foundation, Grants Committee, Washington, USA Australian Research Council, Grants Committee, Canberra, Australia The University of New South Wales, Promotion to Professor, Sydney, Australia Deutsche Forschungsgemeinschaft (DFG), Grants Committee, Germany. Israel Science Foundation (ISF), Grants Committee, Israel. State Scientific Committee (KBN), Grants Committee, Poland. National Scientific Committee (NCN), Grants Committee, Poland.

9 HONOURS

Scientific Secretary, XVIIth Symposium on Advanced Problems and Methods in Fluid Mechanics, Sobieszewo (Poland), Sept.1985. Scientific Secretary, Publication Subcommittee, IUTAM Symposium on Adiabatic Waves in Liquid-Vapour Systems, Göttingen (Germany), Aug. 1989. Invited paper at 2nd European Fluid Mechanics Conference, Warsaw, September 1994 Invited paper at Euromech Colloquium 355 on Interfacial Instabilities, Paris, September 1996 The Polish Academy of Sciences Award (Nagroda Wydzialu IV PAN) for research on jets and droplets, December 1996. International Symposium on Advances in Computational Heat Transfer CHT97, Cesme 1997, memeber of the Scientific Committee and the expository review lecture. European Science Foundation project on Applied Mathematics for Industrial Flow Problems, Polish memeber of the Steering Committee, memeber of the Scientific Committee, “International AMIF conferences”, 1998, 2000, 2002 International Conference on Computational Heat and Mass Transfer, Cyprus, April 1999, memeber of the Scientific Committee. Invited paper at FLUVISU 99, 8me colloque nationale de visualisation et de traitment d’images en mecanique des fluides, Toulose, June 1999. Chairman of the Euromech Colloquium 406 “Image processing methods in applied mechanics”, Warsaw, May 1999. Chairman of the European Science Foundation AMIF Workshop PCC99 “Phase Change with Convection”, Warsaw, June 1999. Member of the Committee of Mechanics of the Polish Academy of Sciences, 1999 - present Memeber of the International Board “Int. Flow Visualization Symposium”, 2000 - present Memeber of the Scientific Committee, “International Symposium on Advances in Computational Heat Transfer CHT01”, Australia 2001. Memeber of the Scientific Committee and invited paper at “International Symposium on Imaging and Visualization in Transport Phenomena VIM-01”, Antalya, 2002. Silver Cross Award by President of Poland for scienitifc achievements, 2002 Coordinator of “Phase Change with Convection”, CISM Advanced School, Udine, 2002. Secretary General, “International Congress of Theoretical and Applied Mechanics”, Warsaw, 2004. Member of the XCCC IUTAM, 2004 - 2012 Member of the EFMCC Euromech, 2007 - 2012 Member of the Scientific Council of the International Centre for Mechanical Sciences, Udine (Italy), 2009 - Member of ECCOMAS Computational Fluid Dynamics Section, 2010 - Chair of the Fluid Mechanics Division of the Committee of Mechanics PAN, 2007 - 2015

10 PRESENTATIONS AT INTERNATIONAL CONFERENCES

Sept. 1977 XIII Biennale on Adv. Fluid Mechanics  Kortowo, Poland
Sept. 1980 Flow Visualization II  Bochum, Germany
Sept. 1981 XV Biennale on Adv. Fluid Mechanics  Jahranka, Poland
Aug. 1984 16th ICTAM Congress  Lyngby, Denmark
April 1985 Euromech Colloquium  Cambridge, UK
April 1986 Taylor’s 100 Birth. Aniv. Symp.  Cambridge, UK
Aug. 1986 Flow Visualization IV  Paris, France
Nov. 1987 ICALEO  San Diego, USA
April 1988 GAMM  Wien, Austria
Aug. 1988 17th ICTAM Congress  Grenoble, France
Aug. 1989 IUTAM Symposium on Topol. Fluid. Mech.  Cambridge, UK
Sept. 1989 XIX Biennial on Adv. Fluid Mechanics  Kozubnik, Poland
Dec. 1989 10 Aust. Fluid Mech. Conf.  Melbourne, Australia
Aug. 1991 1. Euromech Fluid Mech. Conf.  Cambridge, UK
Dec. 1991 11. ABCM Mech. Eng. Conf.  San Paulo, Brazil
Aug. 1992 18th ICTAM Congress  Haifa, Israel
Oct. 1992 Flow Visualization VI  Yokohama, Japan
Oct. 1993 1. Japan-Polish Joint Seminar  Tokyo, Japan
June 1994 2. Japan-Polish Joint Seminar  Pułtusk, Poland
Aug. 1994 10th Int. Heat Transfer Conference  Brighton, UK
Sept. 1994 2. Euromech Fluid Mech. Conf. (invited)  Warszawa, Poland
Sept. 1994 Jap.-Centr. Europe Workshop  Pułtusk, Poland
June 1995 Euromech Colloquium 335  Rome, Italy
July 1995 3 ICIAM Congress  Hamburg, Germany
Aug. 1996 19th ICTAM Congress  Kyoto, Japan
Sept. 1996 Euromech Colloquium 355 (invited)  Paris, France
Dec. 1996 Heat’96  Kielce, Poland
May 1997 Advances in Computational Heat Transfer (invited) Cesme, Turkey
Sept. 1997 3. Euromech Fluid Mech. Conf.  Göttingen , Germany
June 1998 Advanced Computational Methods in Heat Transfer Cracow , Poland
Sept. 1998 8th Flow Visualization  Sorrento, Italy
Oct. 1998 AMIF - ESF Conference  San Feliu, Spain
May 1999 Euromech 406 Colloquium (chair)  Warszawa, Poland
June 1999 FLUVISU 99 (invited)  Toulouse, France
June 1999 PCC99 AMIF-ESF Workshop (chair)  Warszawa, Poland
Aug. 1999 Cold-Region Thermal Eng. ISTESCR’99  Darmstadt, Germany
Sept. 1999 Laser Anemometry Adv. EALA’99  Rome, Italy

Aug. 2000 9th Flow Visualization  Edinburg, Scotland
Aug. 2000 20th ICTAM Congress  Chicago, USA
June 2001 CMEM 2001  Alicante, Spain
Dec. 2001 3d ISEH  Tempe (AZ), USA
May 2002 VIM’01  Belek, Turkey
June 2002 HEAT2002  Baranow, Poland
Aug. 2002 12 Heat Transfer Conf.  Grenoble, France
Oct. 2002 Eurotherm 71  Reims, France
June 2003 Eurotherm 69 (invited)  Bistra, Slovenia
Aug. 2003 5. Euromech Fluid Mech. Conf.  Toulouse, France
Aug. 2004 21st ICTAM Congress (GenSec)  Warsaw, Poland
May 2005 4th ICCHMT (invited)  Paris-Cachan, France
Sept. 2005 CDMM2005  Warsaw, Poland
June 2006 4th ICNMM ASME  Limerick, Irleand
June 2006 6th Euromech Fluid Mech. Conf.  Stockholm, Sweden
Sept. 2006 12th Flow Visualization  Goettingen, Germany
May 2008 IUTAM Symp. Modelling Nanomat. and Nanosys. (invited) Aalborg, Danmark
June 2008 6th ICNMM ASME  Darmstadt, Germany
Aug. 2008 22nd ICTAM Congres  Adelaide, Australia
Sept. 2008 7th Euromech Fluid Mech. Conf.  Manchester, UK
Feb. 2009 GAMM 2009  Gdansk, Poland
June 2009 ExHFT-7  Krakow, Poland
May 2011 CMM-2011 (invited)  Warsaw, Poland
Sept. 2011 ETC-13  Warsaw, Poland
Aug. 2012 23rd ICTAM Congress  Beijing, China
Sept. 2012 NHT-2012 (invited)  Wroclaw, Poland
Sept. 2012 ECCOMAS2012 (invited)  Wien, Austria
Sept. 2015 PCM-CMM2015 (invited)  Gdansk, Poland
June 2016 17th ISFV (invited)  Gatlinburg, USA
Sept 2016 KKMP2016 (invited)  Belchatow, Poland
Sept. 2018 12th Euromech Fluid Mech. Conf.  Vienna, Austria

11 PATENTS

1. Method of nanofibres production, B. Laszkiewicz, T. Kowalewski, P. Kulpinski, B. Niekraszewicz, P. Czarnecki, M. Rubacha, B. Peczek, J. Jedrzejczak, Patent PL 197202, 31.03.2008 WUP 03/08.

2. Method of producing nanomats, T. A. Kowalewski, D. Lamparska, K. Zembrzycki Patent PL 222733B1, 31.08.2016 WUP 08/16.

3. Application of insulation by nanofibreous polymer mats made with an electrospinning technology for the manufacture of neuroprotecting bandages for use in the prevention of trauma changes in the brain, Andrychowski J., Czernicki Z., Frontczak-Baniewicz M., Golabek-Sulejczak D., Kowalczyk T., Kowalewski T.A., Nakielski P., Patent PL 225858B1, 31.05.2017, WUP 05/17.

4. Use of wire insulations made from polymer nanofibers to prevent scarring after neurosurgery Andrychowski J., Czernicki Z., Frontczak-Baniewicz M., Golabek-Sulejczak D., Kowalczyk T., Kowalewski T.A., Patent PL 226479B1, 31.07.2017, WUP 07/17.

12 PUBLICATIONS

1. Experimental determination of filtration coefficients along bunch of fibres (in Polish), T.A. Kowalewski, W. Kalita, IFTR Reports - Prace IPPT, vol. 28/1976, Warszawa 1976.

2. Optical metod for measuring concentration profiles of droplet suspension flowing through a tube (in Polish), T.A. Kowalewski, IFTR Reports - Prace IPPT, vol. 49/1976, Warszawa 1976.

3. Velocity profiles of suspension flowing through a tube, T.A. Kowalewski, Arch. Mech., 32, pp. 857-865, 1980.

4. Laser Doppler measurements of the velocity profiles of concentrated droplet suspensions flowing through a tube, T.A. Kowalewski, Bericht 118/1980, Max-Planck-Institut für Strömungsforschung, Göttingen 1980.

5. Experimental investigation of laminar flow of suspension in channels (in Polish), T.A. Kowalewski, Ph.D. dissertation, IPPT PAN, Warsaw 1982.

6. Laser method of particles distribution measurements in the flow of suspensions, T.A. Kowalewski, in: Flow Visualization II, Bochum 1980, Ed. W. Merzkirch, Hemisphere 1982, pp. 541-546.

7. Concentration and velocity measurements in the flow of droplet suspensions, T.A. Kowalewski, Exp. in Fluids, 2, pp. 213-219, 1984.

8. An experimental study of the lateral migration of a droplet in a creeping flow, W. Hiller, T.A. Kowalewski, Exp.in Fluids, 5, pp. 43-48, 1987.

9. Simultaneous measurement of temperature and velocity fields in thermal convective flows, W. Hiller, T.A. Kowalewski, in: Flow Visualization IV, Paris 1986, Ed. Claude Veret, Hemisphere 1987, pp. 617-622.

10. Eine einfache Hochgeschwindigkeitskamera mit CCD- Sensor, W. Hiller, T.A. Kowalewski, Bericht 8/1987, Max-Planck- Institut für Strömungsforschung, Göttingen 1987.

11. Freistrahlversuche zum genauen Einsatz der Verdampfungseffekte bei Einspritzung von verschiedenen motorischen Kraftstoffen, H. Chaves, T.A. Kowalewski, H.-D. Speckmann and G.E.A. Meier, Bericht 102/1987, Max-Planck-Institut für Strömungsforschung, Göttingen 1987.

12. Untersuchung des Verdampfungsverhaltens von Einspritzstrahlen in einem Saugrohrmodell, H. Chaves, T.A. Kowalewski, and G.E.A. Meier, Bericht 103/1987, Max-Planck-Institut für Strömungsforschung, Göttingen 1987.

13. An optical method for surface tension measurements of dispersed liquid droplets, W.J. Hiller, T.A. Kowalewski, in: Proc. of ICALEO’87, 6th Int. Congress on Application of Laser and Electro-Optics, Optical Methods in Flow and Particle Diagnostics, vol. 63. Ed. W. Stevenson, Publ. Laser Institut of America, pp. 106-110, Toledo 1988.

14. Three-dimensional structures in laminar natural convection in a cubic enclosure, W.J. Hiller, T.A. Kowalewski, St. Koch, Proc. of 1st. Word Conference on Exp. Heat Transfer, Fluid Mech., and Thermodyn., Dubrovnik 1988, ed. R.K. Shah et al., Elsevier Scis. Publ., New York, Amsterdam, London, pp. 722-729, 1988.

15. Similarity in the behaviour of initially saturated or subcooled liquid jets discharging through a nozzle, H. Chaves, T.A. Kowalewski, Th. Kurschat, G.E.A. Meier, E.A. Müller, Chem. Phys. 126, pp. 137-143, 1988.

16. Simultane Erfassung von Temperatur- und Geschwindigkeitsfeldern in einer thermischen Konvektionsströmung mit ungekapselten Flüssigkristalltracern, W.J. Hiller, T.A. Kowalewski, St. Koch, in: 2D-Messtechnik DGLR-Workshop 18-19. Oct. 1988, Markdorf, DGLR-Bericht 88-04, pp. 31 - 39, Bonn 1988.

17. Application of the frame transfer charge-coupled device for high speed imaging, W.J. Hiller, T.A. Kowalewski, Optical Eng. 28 pp. 197-200, 1989; also in: Proc. 18th Int. Congr. of High Speed Photography and Photonics, 28. Aug.-2. Sept. 1988 Xian China, SPIE Vol. 1032, pp. 763 - 765, Washington 1990.

18. Three-dimensional structures in laminar natural convection in a cubic enclosure, W.J. Hiller, T.A. Kowalewski, St. Koch, Exp. Thermal and Fluid Sci., 2, pp. 34-44, 1989.

19. Surface tension measurements by the oscillating droplet method, W.J. Hiller, T.A. Kowalewski, Physico-Chem. Hydrodynam., 11, pp. 103-112, 1989.

20. Optical investigation of oscillating liquid droplets, W.J. Hiller, T.A. Kowalewski, ZAMM 69(6), pp. 629-630, 1989.

21. Schnelle Bildaufzeichnung mit CCD-Kameras und gepulsten LEDs - High Speed Image Recording, W.J. Hiller, T.A. Kowalewski, B. Stasicki, Laser und Optoelektronic 21, pp. 64-67, 1989.

22. Experimental analysis of free oscillating liquid drops, W.J. Hiller, T.A. Kowalewski, in: Proceedings of the 10th Australasian Fluid Mechanics Conference, University of Melbourne, Dec. 1989, pp. 7.21 -7.24, 1989.

23. Liquid microjets - a useful tool for the measurement of material properties, W.J. Hiller, T.A. Kowalewski, in: Proc. of the 10th ABCM Mechanical Engineering Conf., Rio de Janeiro (Brasil), Dec. 1989, Eds. M. Hirata et al., pp. 423 - 426, COPPE/UFRJ 1989.

24. Vergleichende Grunlagenstudie über das Schwingungsverhalten osczillierender Wurzelkanalinstrumente, Griesinger H.R., W.J. Hiller, T.A. Kowalewski, Kreter F., Zahnärztliche Praxis 5, pp. 168-173, 1989.

25. Vibrationsinduzierte Strömungsfelder und Kavitationseffekte dreier Systeme zur Wurzelkanalaufbereitung im Modellversuch, Griesinger H.R., W.J. Hiller, T.A. Kowalewski, Kreter F., Zahnärztliche Praxis 6, pp. 213-217, 1989.

26. Experimental and Numerical investigation of natural convection in a cube with two heated side walls, W.J. Hiller, T.A. Kowalewski, St. Koch, G. de Vahl Davis, M. Behnia, in: Proc. of IUTAM symposium on topological fluid mechanics, Cambridge U.K., Aug. 13-18, 1989, pp. 717 - 726, CUP 1990.

27. Experimental and theoretical investigations of large amplitude oscillations of liquid droplets, E. Becker, W.J. Hiller, T.A. Kowalewski, J. Fluid Mech. 231, pp. 189-210, 1991

28. Unsteady droplet evaporation W.J. Hiller, T.A. Kowalewski, in: Proc. of 11th ABCM Mech. Eng. Conf. São Paulo (Brasil), Dec. 1991, Ed. Esp. da Revista Brasileira de Ciências Mecânicas São Paulo, pp. 17-20, 1991.

29. Visualization of 3-D natural convection - comparison with numerical results, W.J. Hiller, T.A. Kowalewski, St. Koch, K. Range, M. Behnia, G. de Vahl Davis, in: Proc. of 11th ABCM Mech. Eng. Conf. São Paulo (Brasil), Dec. 1991, Ed. Esp. da Revista Brasileira de Ciências Mecânicas São Paulo, pp. 21-24, 1991.

30. Computer-aided discrimination of slow and fast tracer paths, H.-H. Bartels-Lehnhoff, W.J. Hiller, T.A. Kowalewski, Exp. in Fluids, 13, pp. 239-248, 1992; also in: Proceedings of The Sixth International Symposium on Flow Visualisation, Yokohama 1992, Eds. Tanida Y. & Miyashiro H., Springer-Verlag, pp. 873–877, 1992.

31. Visualisation of 3-D natural convection, W.J. Hiller, T.A. Kowalewski, St. Koch, P. Mitgau, K. Range, Proceedings of The Sixth International Symposium on Flow Visualisation, Yokohama 1992 Eds. Tanida Y. & Miyashiro H., Springer-Verlag, pp. 674–678, 1992.

32. Charge-coupled devices in flow visualisation, W.J. Hiller, T.A. Kowalewski, V. Llorach Forner, B. Stückrad, M. Behnia, in: Proceedings of The Sixth International Symposium on Flow Visualisation, Yokohama 1992, Eds. Tanida Y. & Miyashiro H., Springer-Verlag, pp. 695–699, 1992.

33. Behaviour of small diameter evaporating jets, T.A. Kowalewski, W.J. Hiller, Behnia M.,, in: Proceedings of the 11th Australasian Fluid Mechanics Conference, University of Tasmania, Hobart, Australia, 14-18. Dec. 1992, pp. 905–908, U.T. 1992.

34. High speed frame transfer CCD, W.J. Hiller, T.A. Kowalewski, Tatarczyk Th., in: Proc. 20th Int. Congr. of High Speed Photography and Photonics, 21-25. Sept. 1992 Victoria, Canada, SPIE Vol. 1801, pp. 595–601, Washington 1993.

35. Wake patterns of a piston gliding in a rotating circular duct, W.J. Hiller, T.A. Kowalewski, in: Bluff-Body Wakes, Dynamics and Instabilities, IUTAM Symposium, Göttingen, Germany, Sept. 7-11, 1992, Eds. H. Eckelmann et al., Springer-Verlag, pp. 165–169, 1993.

36. Onset of natural convection in a cube, W.J. Hiller, St. Koch, F. Stella, T.A. Kowalewski, Int. J. Heat Mass Transfer, 13, pp. 3251–3263, 1993.

37. An experimental study of evaporating small diameter jets, T.A. Kowalewski, W.J. Hiller, M. Behnia, Physics of Fluids A 5, pp. 1883–1890, 1993.

38. Measurement of dynamic surface tension by the oscillating droplet method, B. Stückrad, W.J. Hiller, T.A. Kowalewski, Exp. in Fluids, 15, pp. 332-340, 1993.

39. Nonlinear dynamics of viscous droplets, E. Becker, W.J. Hiller, T.A. Kowalewski, J. Fluid Mech., 258, pp. 191-216, 1994.

40. Computational and experimental visualisation in heat and mass transfer problems, T.A. Kowalewski, W.J. Hiller and G. de Vahl Davis, in Proc. of the First Japanese-Polish Joint Seminar in Advanced Computer Simulation, Tokyo, Nov. 8-9, 1993, pp. 60–69, Edts. Akiyama, Kleiber, Wolański, University of Tokyo Feb. 1994.

41. Verfolgung von Teilchen in einer dreidimensionalen Strömung, T.A. Kowalewski, P.M. Mitgau, W.J. Hiller, ZAMM 74(5), pp. T394–396, 1994.

42. Reduction of nonlinear dynamic systems by phase space analysis, T.A. Kowalewski, E. Becker, U. Brosa, Computer Assisted Mechanics and Eng. Sci., 1, pp. 39-48, 1994.

43. Experimental and numerical study of three-dimensional natural convection and freezing in water, C. Abegg, T.A. Kowalewski, G. de Vahl Davis, W.J. Hiller, St. Koch, E. Leonardi, G.H. Yeoh, Proc. of 10th International Heat Transfer Conference, Brighton, England, Edt. G.F. Hewitt, vol.4, pp. 1-6, IChemE 1994.

44. Nonlinear Oscillations of Viscous Droplets, T.A. Kowalewski, D. Bruhn, Proc. of Japanese-Centr. European Workshop on Adv. Comp. in Eng., Pułtusk 1994, pp. 63-68, Edts. Akiyama, Kleiber, IPPT PAN Warszawa 1994.

45. Experimental research on briquette destruction caused by refraction waves, M. Gawor, J. Rysz, A. Smolarski, T.A. Kowalewski, Archives of Mining Scs., vol. 39, pp. 313-330, 1994.

46. Selected free surface flow problems - liquid jets and drops, T.A. Kowalewski, (in Polish) Habilitation, Prace IPPT PAN 3/1995, pp.1-98, Warszawa 1995.

47. Experimental investigations of oscillating liquid droplet, T.A. Kowalewski, D. Bruhn, W.J. Hiller, F. Obermeier, ZAMM vol. 76/5, pp.77-78, 1996.

48. On the separation of droplets from a liquid jet, T.A. Kowalewski, Fluid Dynamics Research 17, pp. 121-145, 1996.

49. Distribution of particles suspended in 3-d laminar convection flow, A. Yarin, T.A. Kowalewski, W.J. Hiller, St. Koch, Physics of Fluids 8, pp. 1130-1140, 1996.

50. Experimental and numerical investigations of natural convection in freezing water, T.A. Kowalewski, A. Cybulski, in Int. Conf. on Heat Transfer with Change of Phase, Mechanics vol. 61 p.II, pp. 7-16, Kielce 1996.

51. Experimental validation of numerical codes in thermally driven flows, T.A. Kowalewski, Prace IPPT PAN 4/1997, pp.1-20, Warszawa 1997.

52. Natural convection with phase change (in Polish), T.A. Kowalewski, A. Cybulski, Prace IPPT PAN 8/1997, pp.1-58, Warszawa 1997.

53. Estimation of acoustic streaming: theoretical model, Doppler measurements and optical visualisation, A. Nowicki, T.A. Kowalewski, W. Secomski, J. Wojcik, European J. of Ultrasound 7, pp. 73-81, 1998.

54. Experimental validation of numerical codes in thermally driven flows, T.A. Kowalewski, in Adv. in Computational Heat Transfer, G. de Vahl Davis, E. Leonardi (eds), Begel House Inc., pp.1-15, New York 1998.

55. An experimental benchmark for freezing water in the cubic cavity T.A. Kowalewski, M. Rebow, in Adv. in Computational Heat Transfer, G. de Vahl Davis, E. Leonardi (eds), Begel House Inc., pp.149-156, New York 1998.

56. Fixed grid finite elemnt analysis of solidification J. Banaszek, T.A. Kowalewski, M. Rebow, in Adv. in Computational Heat Transfer, G. de Vahl Davis, E. Leonardi (eds), Begel House Inc., pp.471-478, New York 1998.

57. Free surface natural convection in a differentially heated rectangular cavity, T.A. Kowalewski, G. de Vahl Davis, E. Leonardi, in Advanced Computational Methods in Heat Transfer V, A.J. Nowak, C.A. Brebbia, R. Bialecki, M. Zerroukat (eds.), pp. 145-154, CMP 1998.

58. Particle Image Velocimetry with Optical Flow, G. M. Quénot, J. Pakleza, T.A. Kowalewski, Experiments in Fluids 25, pp.177-189, 1998.

59. Analysis of validation of the semi-implicit FEM algorithm for natural convection and freezing water, J. Banszek, T.A. Kowalewski, M. Rebow, K. Blogowska, in Computational Technologies for Fluid/Thermal/Structural/Chemical Systems with Industrial Applications, V.V. Kudriavtsev, W. Cheng (eds), ASME PVP-Vol.377-1, pp.119-126, 1998.

60. Fluid Image Velocimetry for Unseeded flow, Y. Levy, B. Golovanevsky T.A. Kowalewski, in Proc. of 9th Int. Symp. on Application of Laser Technique in Fluid Mech., July 13-16, Lisbon, vol. II, pp. 34.7.1-34.7.9, 1998.

61. Characterisation of the interaction between a boundary layer and a cavity using Digital Particle Velocimetry with Optical Flow, G. Quenot, A. Rambert, P. Gougat, T.A. Kowalewski, in Proc. of 9th Int. Symp. on Application of Laser Technique in Fluid Mech., July 13-16, Lisbon, vol. II, pp.39.4.1-39.4.7, 1998.

62. Study of free convection in binary systems with phase change, T.A. Kowalewski, J. Banaszek, M. Rebow, P. Furmanski, T. Wisniewski, in Heat Transfer 1998, Kyongju, Korea, edt. J.S. Lee, vol.7, pp. 211-216, Taylor and Francis Inc., 1998.

63. Particle Image Velocimetry and Thermometry in Freezing Water, T.A. Kowalewski, A. Cybulski, M. Rebow, in 8th Int. Symposium on Flow Visualization, Sorrento, eds. G.M. Carlomagno and I. Grant, CD ROM Proceedings ISBN 0953399109 , pp. 24.1-24.8, Edinburgh, 1998.

64. Particle Image Velocimetry using Optical Flow for Image Analysis, T.A. Kowalewski, G. Quenote, J. Pakleza, in 8th Int. Symposium on Flow Visualization, Sorrento, eds. G.M. Carlomagno and I. Grant, CD ROM Proceedings ISBN 0953399109, pp. 47.1-47.11, Edinburgh, 1998.

65. Infrared and Liquid Crystal Thermography in Natural Convection, T.A. Kowalewski, T. Wisniewski, M. Rebow, K. Blogowska, in 8th Int. Symposium on Flow Visualization, Sorrento, eds. G.M. Carlomagno and I. Grant, CD ROM Proceedings ISBN 0953399109, pp. 212.1-212.8, Edinburgh, 1998.

66. Thermochromic liquid crystals in heat transfer research, T.A. Kowalewski,, J. Stasiek, invited paper at Heat Mass Transfer Conference, Swieradow, Sept. 1998, Prace Polit. Wroclaw. 53/9, vol. 1, pp. 401-424, Wroclaw 1998.

67. Freezing of water in the differentially heated cubic cavity, T.A. Kowalewski, M. Rebow, Int. J. of Comp. Fluid Dyn. vol. 11, pp. 193-210, 1999.

68. Particle Image Velocimetry and Thermometry in Freezing Water, T.A. Kowalewski, A. Cybulski, frontispiece (colour page) in J. of Visualization, vol. 1, p. 335 , 1999.

69. Effect of finite wall conductivity on flow structures in natural convection, E. Leonardi, T.A. Kowalewski, V. Timchenko, G. de Vahl Davis, paper at CHMT99, Proceedings of the International Conference on Computational Heat and Mass Transfer, Cyprus 26-29 April, 1999, eds. A.A. Mohamad & I. Sezai, pp. 182-188, Eastern Mediterranean University Printinghouse 1999.

70. Particle Image Velocimetry and Thermometry Using Liquid Crystals, T.A. Kowalewski, invited paper at FLUVISU 99, 8me colloque nationale de visualisation et de traitment d’images en mecanique des fluides, Toulose, June 1-4, 1999, ENSICA, pp. 33-48, Image Service 1999.

71. Measurement of Temperature and Velocity of Freezing Water using Liquid Crystal Tracers T.A. Kowalewski, in Advances in Cold-Region Thermal Engineering and Sciences, eds. K. Hutter, Y. Wang and H. Beer, Lecture Notes in Physics vol. 533, pp. 83-90, Springer Verlag, Berlin 1999.

72. Particle Image Velocimetry for Vapour Bubble Growth Analysis, T.A. Kowalewski, J. Pakleza, A. Cybulski; Proceedings of the 8th International Conference on Laser Anemometry, pp. 243-250, Univ. of Rome La Sapienza, 1999.

73. An axisymmetry-breaking instability of axially symmetric natural convection A.Yu. Gelfgat, P.Z. Bar-Yoseph, A. Solan, T.A. Kowalewski, Int. J. Trans. Phenomena: vol.1, pp. 173-190, 1999.

74. Semi-implicit FEM analysis of natural convection in freezing water J. Banaszek, Y. Jaluria, T.A. Kowalewski, M. Rebow; Num. Heat Transfer, Part A, vol. 36, pp. 449-472, 1999.

75. Phase-change problems with free convection: fixed grid simulation M. Giangi, T.A. Kowalewski, F. Stella; Comp. & Vis. in Scs. vol.2 (2/3), pp. 123-130, 1999.

76. Large-scale structures forming in a cross flow: particle image velocimetry conditional analysis R. Camussi, A. Stella, T.A. Kowalewski, G. Guj, Machine Graphics & Vision vol. 8 (4), pp. 579-595, 1999.

77. Natural convection during ice formation: numerical simulation vs. experimental results M. Giangi, T.A. Kowalewski, F. Stella, E. Leonardi, CAMES, vol 7(3), pp. 321-342, 2000.

78. Natural convection with phase change in one-component and binary systems (in Polish), J. Banaszek, T.A. Kowalewski, P. Furmanski, M. Rebow, A. Cybulski, T.S. Wisniewski; Prace IPPT PAN 3/2000, pp.1-132, Warszawa 2000.

79. Tracking of thermal structures from infrared camera by PIV method, T.A. Kowalewski, G. Hetsroni, B. Hu, A. Mosyak in 9th Int. Symposium on Flow Visualization, Edinburgh, Edt. I. Grant, CD ROM Proceedings, pp. 131.1-9, Edinburgh, 2000.

80. Visualization of vapour bubble growth, T.A. Kowalewski, J. Pakleza, J.-B. Chalfen, M.-C. Duluc, A. Cybulski in 9th Int. Symposium on Flow Visualization, Edinburgh, Edt. I. Grant, CD ROM Proceedings, pp. 176.1-9, Edinburgh, 2000.

81. Tracking of coherent thermal structures on a heated wall by means of IR thermography T.A. Kowalewski, G. Hetsroni, B. Hu, A. Mosyak; Experiments in Fluids, vol. 30, pp. 286-294, 2001.

82. Experimental model for casting problems, T.A. Kowalewski, A. Cybulski, T. Sobiecki, in Computational Methods and Experimental Measurements, Eds. Y. V. Esteve, G.M. Carlomagno, C.A. Brebia, pp. 179-188, WIT Press, Southampton 2001.

83. Experimental investigations on thermal, thermocapillary, and forced convection in Czochralski crystal growth configuration, J. Aleksic, J.A. Szymczyk, A. Leder, T.A. Kowalewski, in Computational Methods and Experimental Measurements, Eds. Y. V. Esteve, G.M. Carlomagno, C.A. Brebia, pp. 627-636, WIT Press, Southampton 2001.

84. Visualization of Vapour Bubble Growth (fontipiece), T.A. Kowalewski, J. Pakleza, J.-B. Chalfen, M.-C. Duluc, A. Cybulski, J. of Visualization, vol. 4, p. 212, 2001

85. Methodologie experimentale pour l’etude de la croissance de bulles de vapour en site isole, J. Pakleza, T.A. Kowalewski, J.-B. Chalfen, M.-C. Duluc, A. Cybulski, in Science et technologies thermiques face au XXIe siecle, SFT2001, Acte du congres annuel Nantes, Edts. C. Castelain, D. Delaunay, pp. 467-472, Elsevier 2001,

86. Particle Image Velocimetry and Thermometry using Liquid Crystals Tracers, T.A. Kowalewski, PIV’01 4th Int. Symp. on Particle Image Velocimetry CD ROM Proc, pp.1134.1-10, Edt. J. Kompemhans, DLR Mitteilung 2001-03, Goettingen 2001.

87. Vertical transport ond mixing in the Salt-Lake basin, T.A. Kowalewski, P. Monti, M. Princevac, W.C. Chan, H.J. Fernando, 3d Int. Symp. on Enviromental Hydraulics, CD ROM Proc, pp.00286.1-6, Edts. D. Boyer & R. Rankin, ASU, Tempe 2001.

88. Slope flow measurements during vertical transport and mixing (VTMX) field experiment, Salt Lake City 2000, M. Princevac, T.A. Kowalewski, H.J. Fernando, P. Monti, W.C. Chan, J. Anderson, 3d Int. Symp. on Enviromental Hydraulics, CD ROM Proc, pp.00238.1-6, Edts. D. Boyer & R. Rankin, ASU, Tempe 2001.

89. Application of liquid crystal tracers for full field temperature and velocity measurements, T.A. Kowalewski, 3d Int. Symp. on Enviromental Hydraulics, CD ROM Proc, pp.00011.1-6, Edts. D. Boyer & R. Rankin, ASU, Tempe 2001.

90. Thermochromic liquid crystals applied for heat transfer research, J.A. Stasiek, T.A. Kowalewski, Opto-electronics review, vol. 10(1), pp. 1-10, 2002

91. Observations of flow and turbulence in the nocturnal boundary layer, P. Monti, H.J. Fernando, M. Princevac, W.C. Chan, E.R. Padyjak, T.A. Kowalewski, J. Atmospheric Scs. vol. 59(17), 2513-2534, 2002

92. Experimental investigation of vapour bubble growth, T.A. Kowalewski, M.-C. Duluc, J. Pakleza, Heat Transfer 2002, vol. 3, pp. 479-485, Elsevier 2002

93. Experimental benchmark for casting problems, T.A. Kowalewski, A. Cybulski, T. Michalek, Heat Transfer 2002, vol. 4, pp. 813-819, Elsevier 2002

94. Particle Image Velocimetry and Thermometry in two phase problems, T.A. Kowalewski, in Visualization and Imaging in Transport Phenomena, Annals of New York Academy of Sciences, vol.972, pp.213-222, edts. S. Siedeman & A. Landesberg, NY 2002.

95. Tracking of coherent thermal structures on a heated wall. Part 2: DNS Simulation, T.A. Kowalewski, A. Mosyak, G. Hetsroni, Exp. in Fluids 34, pp. 390-396, 2003.

96. Simulation of the water freezing process - numerical benchmarks, T. Michalek, T.A. Kowalewski, TASK Quarterly 7: 389-408, 2003.

97. Experimental Methods for Quantitative Analysis of Thermally Driven Flows, T.A. Kowalewski, in Phase Change with Convection, eds. T. Kowalewski, D. Gobin, CISM Course Lecture Notes, vol.449, pp. 167-215, Springer 2004

98. Experimental analysis of vapour bubble growth on heated surface, T.A. Kowalewski, K. Pakleza, R. Trzcinski, A. Zachara, Archives of Theromdynamics, 25, pp. 1-12, 2004

99. Visualization of the Flow Structure and Temperature Field in the Region of Mixed Convection, .E. Fornalik , W. Leiner, J. S. Szmyd, H. Ozoe, T.A. Kowalewski, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

100. Particle Image Velocimetry (PIV) for Cloud Droplets - Laboratory Investigations, P. M. Korczyk,S. P. Malinowski,T.A. Kowalewski, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

101. Experimental and Numerical Study of Marangoni - Natural Convection, Le Han Tan, E. Leonardi, T. J. Barber, S. S. Leong, T.A. Kowalewski, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

102. AHP Setup for Low Laminar Melt Flow Study in Crystal Growth, M. A. Gonik, J. A. Szymczyk, T.A. Kowalewski, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

103. Electrospinning of Liquid Jets, S. Blonski,T.A. Kowalewski, A. Blasinska, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005

104. Natural Convection for Anomalous Density Variation of Water - Numerical Benchmark, T. Michalek, T.A. Kowalewski, B. Sarler, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

105. Thermodynamic Parameters of Vapour Bubble Growth by Image Analysis, Zachara, J. Pakleza, S. Blonski, R. Trzcinski, T.A. Kowalewski, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

106. Experiments on Up-slope to Down-slope Transition in an Inclined Box Filled with Water, K. Dekajlo,T.A. Kowalewski, H. J. S. Fernando, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

107. Optical technique DPIV in measurements of granular material flows, Part 1/3 - plane hoppers, I. Sielamowicz, T.A. Kowalewski, S. Blonski, Chem. Eng. Sci., 60, pp. 589-598, 2005

108. Laboratory benchmarks for validating numerical simulation of casting problems, T.A. Kowalewski, A. Cybulski, T. Michalek, M. Kowalczyk, Prace IPPT - IFTR Reports, vol. 1/2005, IPPT PAN, Warszawa 2005

109. Natural Convection for Anomalous Density Variation of Water - Numerical Benchmark, T. Michalek, T.A. Kowalewski, B. Saler, Progress in Computational Fluid Dynamics 5, pp. 158– 170, 2005.

110. Numerical Benchmark Based on Natural Convection of Freezing Water , T.A. Kowalewski, T. Michalek, 4th International Conference on Computational Heat and Mass Transfer, vol 1, pp. 7-11, University Cachan 2005

111. Experimental and numerical investigation of natural convection in rectangular inclined enclosure, K. Dekajlo, T.A. Kowalewski, w Numerical Heat Transfer 2005,Eurotherm 82, Sept. 13-16, 2005 Krakow, pp. 953-962.

112. Selected laboratory benchmarks for validating numerical simulation of casting problems, T.A. Kowalewski, M. Kowalczyk, T. Michalek, A. Cybulski w Numerical Heat Transfer 2005, Eurotherm 82, Sept. 13-16, 2005 Krakow, pp. 589-598.

113. Experiments and modelling of electrospinning process, T.A. Kowalewski, S, Blonski, S. Baral, Bulletin Polish Acad. Scs., vol. 53/4, pp. 385-394, 2005.

114. Mixing of cloud and clear air in centimetre scales observed in laboratory, P. Korczyk, T.A. Kowalewski,, S.P. Malinowski, J. Atmospheric Research 82, pp. 173-182, 2006

115. Digital Particle Image Velocimetry (DPIV) technique in measurements of granular material flows, Part 2 of 3-converging hoppers I. Sielamowicz, T.A. Kowalewski, S.Blonski, Chem. Eng. Scs. 61,5307-5317, 2006.

116. Turbulent Flow in a Micro-Channel, T.A. Kowalewski, S. Blonski, P. Korczyk, in ASME 4th International Conference on Nanochannels, Microchannels and Minichannels, June 2006, Limerick, CD-ROM Proceedings ISBN 0-7918-3778-5, 96090, p. 1-8, 2006.

117. Investigations of turbulence statistics in the laboratory model of an atmospheric cloud, T.A. Kowalewski, P. Korczyk, S. Malinowski in 12th Int. Symposium on Flow Visualization, CD ROM Proceedings, ISBN 0-9533991-8-4, pp. 134.1-10, Goettingen, Optimage Ltd 2006.

118. DPIV technique for granular material flows in silo models, T.A. Kowalewski, I. Sielamowicz in 12th Int. Symposium on Flow Visualization, CD ROM Proceedings, ISBN 0-9533991-8-4, pp. 164.1-10, Goettingen, Optimage Ltd 2006.

119. Analysis of turbulence in micro-channel emulsifier, S. Blonski, P. Korczyk, T.A. Kowalewski, International Journal of Thermal Sciences vol. 46, pp. 1123-1141, 2007

120. PIV analysis of turbulent flow in a micro-channel, S. Blonski, T.A. Kowalewski, JTAM vol. 45, pp. 489-503, 2007

121. Thermochromic Liquid Crystals, T.A. Kowalewski, Chapter 22 in Springer Handbook of Experimental Fluid Mechanics, C. Tropea, A.L. Yarin, J.F. Foss (Eds.), pp.487-561, Springer Verlag Berlin Heidelberg 2007, ISBN: 978-3-540-25141-5

122. Metoda pomiaru ruchu malej grupy czastek opadajacej w cieczy, S. Alabrudzinski, A.L. Ekiel-Jezewska, T.A. Kowalewski, W. Suchecki, in Wybrane zagadnienia przeplyww i wymiany ciepla, W. Suchecki (Eds.), chap. 3, pp.59-85, Oficyna Wyd. Polit. Warsz. 2008.

123. Eksperymentalna analiza przeplyww w skali mikro i nano, T.A. Kowalewski, S. Blonski, P. Korczyk, in Wybrane zagadnienia przeplyww i wymiany ciepla, W. Suchecki (Eds.), chap. 6, pp.127-149, Oficyna Wyd. Polit. Warsz. 2008.

124. Electrospinning of bovine serum albumin. Optimization and the use for production of biosensors, T. Kowalczyk A. Nowicka, D. Elbaum, T.A. Kowalewski, Biomacromolecules, 9(7), pp. 2087-2090, 2008.

125. Laboratory and modeling studies of cloud-clear air interfacial mixing: anisotropy of small-scale turbulence due to evaporative cooling, Sz. Malinowski, M. Andrejczuk, W. W. Grabowski, P. Korczyk, P. Smolarkiewicz, T.A. Kowalewski, New Journal of Physics 10, 075020, 2008.

126. Low-Reynolds-Number Instability of the Laminar Flow Between Wavy Walls, T.A. Kowalewski, J. Szumbarski, S. Blonski, in ASME 6th International Conference on Nanochannels, Microchannels and Minichannels, June 2008, Darmstadt, CD-ROM Proceedings ISBN 0-7918-3826-9, 62070, pp. 1-8, ASME 2008.

127. Flow destabilization and chaotic mixing in the channel with transversely corrugated walls, S. Blonski, J. Szumbarski, T.A. Kowalewski, in XXII International Congress of Theoretical and Applied Mechanics, Adelaide, Australia, August 25-29, 2008, CD-ROM proceedings ISBN 978-0-9805142-1-6

128. Evidence of anisotropy of small scale turbulence in the laboratory model of an atmospheric cloud, P. Korczyk, T.A. Kowalewski, Sz. Malinowski, in XXII International Congress of Theoretical and Applied Mechanics, Adelaide, Australia, August 25-29, 2008, CD-ROM proceedings ISBN 978-0-9805142-1-6

129. Nano-scale solidification phenomena, T.A. Kowalewski, J. Czerwinska, E. Deriaz, in XXII International Congress of Theoretical and Applied Mechanics, Australia, August 25-29, 2008, CD-ROM proceedings ISBN 978-0-9805142-1-6

130. Heat transfer enhancement through the use of transverse vortex generators, S. Blonski, T.A. Kowalewski, D. Mikielewicz, A. Stasiek, J. Stasiek, in Proceedings of XXVI Congresso Nazionale UIT sulla Trasmissione del Calore, Palermo, 23-25 Giugno 2008

131. Turbulent mixing of cloud with the environment: Two-phase evaporating flow as seen by particle imaging velocimetry, Sz. P. Malinowski P. Korczyk, T.A. Kowalewski, in Proceedings of International Conference on Clouds and Precipitation ICCP-2008, 7-11 July, 2008 Cancun, Mexico 132. Modeling Electrospinning of Nanofibers, T.A. Kowalewski, S. Barral, T. Kowalczyk, in Mod-

elling Nanomaterials and Nanosystems. R. Pyrz, J.C. Rauhe (Eds.), Proceedings of the IUTAM Symposium, Aalborg, Denmark, 19-22 May, 2008 : IUTAM Bookseries Vol. 13, Springer Science+Business Media B.V., 2009, pp. 279-292.

133. Dynamics of nanofibers conveyed by low Reynolds number flow in a micro-channel, K. Sadlej, E. Wajnryb, M.L. Ekiel-Jezewska, T.A. Kowalewski, in Proceedings of 7th World Conference on Experimental Heat Transfer, Fluid Mechanics and Thermodynamics, 28 June 03 July 2009, Krakow, Poland, J. S. Szmyd, J. Spalek, T. A. Kowalewski (Eds.), CD-ROM, ISBN 978-837464-235-4, AGH 2009, pp. 237-244.

134. Turbulence in micro-channels, S. Blonski, T.A. Kowalewski, in Proceedings of 7th World Conference on Experimental Heat Transfer, Fluid Mechanics and Thermodynamics, 28 June 03 July 2009, Krakow, Poland, J. S. Szmyd, J. Spalek, T. A. Kowalewski (Eds.), CD-ROM, ISBN 978-837464-235-4, AGH 2009, pp. 361-368.

135. Particle clusters settling under gravity in a viscous fluid, S. Alabrudzinski, M. L. Ekiel-Jezewska, D. Chehata-Gomez, T.A. Kowalewski, Physcis of Fluids 21, 073302, 2009.

136. Modelling Electrospinning of Nanofibres, T.A. Kowalewski, S. Barral, PAMM vol. 9, p.463-464, 2009

137. Electrospun nanofibrous biodegradable polyester coatings on Bioglass-based glass-ceramics for tissue engineering, O. Bretcanu, S. K. Misra, D. M. Yunosa, A. R. Boccaccinia, I. Roy, T. Kowalczyk, S. Blonski, T.A. Kowalewski, Materials Chemistry and Physics 118, 420-426, 2009

138. Turbulent mixing of cloud with the enviroment: small scales of two-phase evaporating flow seen by particle imaging velocimetry, Sz. Malinowski, P. Korczyk, T.A. Kowalewski, in Proc. of Euromech Colloquium 512, Torino Oct. 26-29, 2009, eds. D. Tordella, K.S. Sreenivasan, pp. 75-78, Accademia delle Scienze di Torino 2009

139. Empirical description of flow parameters in eccentric flow inside a silo model, I. Sielamowicz, M. Czech, T.A. Kowalewski, Powder Technology 198, pp. 381-394, 2010

140. Dynamics of nanofibres conveyed by low Reynolds number flow in a microchannel, K. Sadlej, E. Wajnryb, M.L. Ekiel-Jezewska, D. Lamparska, T.A. Kowalewski, International Journal of Heat and Fluid Flow 31 pp. 996–1004, 2010

141. Mechanika plynow - Dlaczego tak trudno przewidziec ruch plynu?, S. Drobniak, T.A. Kowalewski, in Polskie i swiatowe osiaganiecia nauki: Nauki Techniczne, tom 2, pp. 389-428, Fundacja in. W. Swietoslawskiego, 2010.

142. Validation problems in computational fluid mechanics, T.A. Kowalewski, Computer Assisted Mechanics and Engineering Sciences 18, 39–52, 2011

143. Empirical description of granular flow inside a model silo with vertical walls, Sielamowicz I., Czech M., Kowalewski T.A., Biosystems Engineering 108, pp. 334–344, 2011

144. Hydro-dynamically modified seeding for micro-PIV, S. Blonski, P. Domagalski, M. Dziubinski, T.A. Kowalewski, Arch. of Mech. 63, pp. 163-182, 2011

145. Empirical analysis of eccentric flow registered by the technique inside a silo model, Sielamowicz I., Czech M., Kowalewski T.A., Powder Technology 212, pp. 38 - 56, 2011.

146. Nanofibres for medical applications at Biocentrum Ochota, Kowalewski T.A., Kowalczyk T., Frontczak-Baniewicz M.M., Golabek-Sulejczak D.A., Andrychowski J., Annual Report 2011, Polish Academy of Sciences, 1, 62-65, 2011

147. Impact of transversely-oriented wall corrugation on hydraulic resistance of a channel flow, Szumbarski J., Blonski S., Kowalewski T.A., The Archive of Mechanical Engineering, LVIII, 4, 441-466, 2011

148. Turbulent mixing of clouds with the environment: Small scale two phase evaporating flow investigated in a laboratory by particle image velocimetry, Korczyk P.M., Kowalewski T.A., Malinowski Sz.P., PhysicaD Nonlinear Phenomena 241, 288-296, 2012

149. Optical birefringence and molecular orientation of electrospun polycaprolactone fibers by polarizinginterference microscopy, Kolbuk D., Sajkiewicz P., Kowalewski T.A., European Polymer Journal 48, 275–283, 2012

150. Magnetic Fe doped ZnO nanofibers obtained by electrospinning, Baranowska-Korczyc A., Reszka A., Sobczak K., Sikora B., Dziawa P., Aleszkiewicz M., Klopotowski L.,Paszkowicz W., Dluzewski P., Kowalski B.J., Kowalewski T.A., Sawicki M., Elbaum D., Fronc K., Journal of Sol-Gel Science and Technology 61 (3), 494–500, 2012

151. Analysis of wall effect on the process of diffusion of nanoparticles in a microchannel, Zembrzycki K., S. Blonski S., Kowalewski T.A., Journal of Physics: Conference Series vol. 392, pp. 012014-0 – 11, 2012

152. Nanofiber nets in prevention of cicatrisation in spinal procedures. Experimental study, Andrychowski J, Frontczak-Baniewicz M, Sulejczak D, Kowalczyk T, Chmielewski T, Czernicki Z, Kowalewski T.A., Folia neuropathologica 51 (2), 147–157, 2013.

153. Electrospun nanofiber mat as a protector against the consequences of brain injury, Sulejczak D. , Andrychowski J., Kowalczyk T., Nakielski P. , Frontczak-Baniewicz M., Kowalewski T.A., Folia Neuropathologica 52(1), 56-69, 2014.

154. The Use of Liquid Crystal Thermography in Selected Technical and Medical Applications, Stasiek J., Jewartowski M.. Kowalewski T.A., Journal of Crystallization Process and Technology 4, 46– 59, 2014.

155. Electrospinning of poly(lactic acid)/polyhedral oligomeric silsesquioxane nanocomposites and their potential in chondrogenic tissue regeneration, C. Gomez-Sancheza, T. Kowalczyk, G. Ruiz De Eguino, A. Lopez-Arraiza, A. Infante, C.I. Rodriguez, T.A. Kowalewski, M. Sarrionandia, J. Aurrekoetxea, J. of Biomaterials Science, Polymer Edition 25(8), pp. 802–825, 2014.

156. Is the Poly (L- Lactide- Co- Caprolactone) Nanofibrous Membrane Suitable for Urinary Bladder Regeneration? M. Pokrywczynska, A. Jundzill, J. Adamowicz, T. Kowalczyk, K. Warda, M. Rasmus, L. Buchholz, S. Krzyzanowska, P. Nakielski, T. Chmielewski, M. Bodnar, A. Marszalek, R. Debski, M. Frontczak-Baniewicz, G. Mikulowski, M. Nowacki, T. A. Kowalewski, T. Drewa, PLoS ONE 9(8):e105295. doi:10.1371/journal.pone.0105295, 2014.

157. Ureter Regeneration - The Proper Scaffold Has to Be Defined, T. Kloskowski, A. Jundzill, T. Kowalczyk, M. Nowacki, M. Bodnar, A. Marszalek, M. Pokruwczynska, M. Frontczak-Baniewicz, T. A. Kowalewski, P. Chlosta, T. Drewa, PLoS ONE 9(8):e106023. doi:10.1371/journal.pone.0106023, 2014.

158. Nanofluids and Nanofibres, T.A. Kowalewski, chapter in: Nanomechanics Selected problems, Wydawnictwo Politechniki Krakowskiej, Muc A., Chwał M., Garstecki A., Szefer G. (Eds.), pp.31–50, 2014

159. Experimental and numerical evaluation of drug release from nanofiber mats to brain tissue, P. Nakielski, T. Kowalczyk, K. Zembrzycki, T.A. Kowalewski, J. Biomed. Mater. Res. 103B, 282–291, 2015.

160. Comparative analysis of empirical descriptions of eccentric flow in silo model by the linear and nonlinear regressions I. Sielamowicz, A. Czech, T. A. Kowalewski, Powder Technology 270, 393– 410, 2015.

161. Hydrogel nanofilaments via core-shell electrospinning, P. Nakielski, S. Pawlowska, F. Pierini, V. Liwinska, P. Hajduk, K. Zembrzycki, E. Zabost, T.A. Kowalewski, PLoS ONE 10(6):e0129816. doi:10.1371/journal.pone.0129816, 2015.

162. Synthesis of carbon nanotubes by the laser ablation method: Effect of laser wavelength, J. Chrzanowska, J. Hoffman, A. Mazurkiewicz, T.A. Kowalewski, Z. Szymanski, L. Stobinski, Phys. Status Solidi B 252(8), 1860–1867, 2015.

163. Atomic force microscopy combined with and optical tweezers (AFM/OT), F. Pierini, K. Zembrzycki, P. Nakielski, S. Pawlowska, T.A. Kowalewski, Measurement Science and Technology 27(2), 025904, 2016.

164. New Amniotic Membrane Based Biocomposite for Future Application in Reconstructive Urology, J. Adamowicz, M. Pokrywczynska, J. Tworkiewicz, T. Kowalczyk, S. V. van Breda, D. Tyloch, T. Kloskowski, M. Bodnar, J. Skopinska-Wisniewska, A. Marszalek, M. Frontczak-Baniewicz, T. A. Kowalewski, T. Drewa, PLOS ONE 11(1) e0146012, DOI: 10.1371/journal.pone.0146012, 2016.

165. Micro and nano fluid mechanics, chapter in Advances in Mechanics: Theoretical, Computational and Interdisciplinary Issues, T.A. Kowalewski, P. Nakielski, F. Pierini, K. Zembrzycki, S. Pawlowska, CRC Press, Eds.: Michal Kleiber et al., pp.27–34, 2016

166. Electrospun poly(3-hexylthiophene)/ poly(ethylene oxide)/graphene oxide composite nanofibers: effects of graphene oxide reduction, F. Pierini, M. Lanzi, P. Nakielski, S. Pawlowska, K. Zembrzycki, T.A. Kowalewski, Polym. Adv. Technol. 27(11), pp. 1465-1475, 2016. DOI:10.1002/pat.3816

167. Mobility of highly deformable nanofilaments, S. Pawlowska, P. Nakielski, F. Pierini, K. Zembrzycki, T.A. Kowalewski, in 24th International Congress of Theoretical and Applied Mechanics, Montreal, August 21-26, 2016, e-proceedings ISBN NR16-127/2016E-EPUB, pp. 1196-1197, 2016.

168. Particles double layer evaluation by atomic force microscopy - optical tweezers, F. Pierini, P. Nakielski, S. Pawlowska, K. Zembrzycki, T.A. Kowalewski, in 24th International Congress of Theoretical and Applied Mechanics, Montreal, August 21-26, 2016, e-proceedings ISBN NR16-127/2016E-EPUB, pp. 1204-1205, 2016.

169. New experimental tools in microfluidics, T.A. Kowalewski, in XXII Fluid Mechanics Conference, Belchatow, Sept. 11-14, 2016, e-book pp. 33-34, O. Wyd. Polit. Wrocl., Wroclaw 2016.

170. Single-Material Organic Solar Cells Based on Electrospun Fullerene- Grafted Polythiophene Nanofibers, F. Pierini, M. Lanzi, P. Nakielski, S. Pawlowska, O. Urbanek, K. Zembrzycki, T.A. Kowalewski, Macromolecules 50(13), pp. 4972-4981, 2017.

171. Electrospun Polyaniline-Based Composite Nanofibers: Tuning the Electrical Conductivity by Tailoring the Structure of Thiol-Protected Metal Nanoparticles, F. Pierini, M. Lanzi, P. Nakielski, T. A. Kowalewski, Journal of Nanomaterials, vol. 2017, Article ID 6142140, (2017)

172. Lateral migration of electrospun hydrogel nanofilaments in an oscillatory flow, S. Pawlowska, P. Nakielski, F. Pierini, I.K. Piechocka, K. Zembrzycki, T.A. Kowalewski, PLOS ONE 12(11):e0187815, DOI:10.1371/journal.pone.0187815, 2017.

173. Fibrous polymer nanomaterials for biomedical applications and their transport by fluids: an overview, S. Pawlowska, T. A. Kowalewski, and F. Pierini, Soft Matter 14, pp. 8421-8444, 2018.

174. Polymer-Based Nanomaterials for Photothermal Therapy: From Light-Responsive to Multifunctional Nanoplatforms for Synergistically Combined Technologies, F. Pierini, P. Nakielski, O. Urbanek, S. Pawlowska, M. Lanzi, L. De Sio, T. A. Kowalewski, Biomacromolecules 19(11), pp. 4147-4167, 2018.

175. Thermoplasmonic activated hydrogel based dynamic light attenuator, Pierini F., Guglielmelli A., Urbanek O., Nakielski P., Pezzi L., Buda R., Lanzi M., Kowalewski T.A., De Sio L., Advanced Optical Materials, ISSN: 2195-1071, DOI: 10.1002/adom.202000324, pp.2000324-1-7, 2020.

176. Ultraviolet light assisted electrospinning of core shell fully cross-linked P(NIPAAm–co–NIPMAAm) hydrogel-based nanofibers for thermally induced drug delivery self regulation , Pawlowska S., Rinoldi C., Nakielski P., Ziai Y., Urbanek O., Li X., Kowalewski T.A., Ding B., Pierini F., Advanced Materials Interfaces, ISSN: 2196-7350, DOI: 10.1002/admi.202000247, pp.2000247-1-13, 2020.

177. Multifunctional platform based on electrospun nanofibers and plasmonic hydrogel: A smart nanostructured pillow for near-infrared light-driven biomedical applications, Nakielski P, Pawowska S, Rinoldi C, Ziai Y, De Sio L, Urbanek O, Zembrzycki K, Pruchniewski M, Lanzi M, Salatelli E, Calogero A, Kowalewski TA, Yarin AL, Pierini F, ACS Applied Materials and Interfaces, ISSN: 1944-8244, DOI: 10.1021/acsami.0c13266, pp.1-15, 2020.

178. Dynamics of electrospun hydrogel filaments in oscillatory microchannel flows: a theoretical and experimental approach, Sankaran A., Pawlowska S., Pierini F., Kowalewski T.A., Yarin A.L., PHYSICS OF FLUIDS, ISSN: 1070-6631, DOI: 10.1063/5.0011005, Vol.32, No.7, pp.072008-1-13, 2020

179. Three-dimensional printable conductive semi-interpenetrating polymer network hydrogel for neural tissue applications, Rinoldi C., Lanzi M., Fiorelli R., Nakielski P., Zembrzycki K., Kowalewski T., Urbanek O., Grippo V., Jezierska-Wozniak K., Maksymowicz W., Camposeo A., Bilewicz R., Pisignano D., Sanai N., Pierini F., Biomacromolecules, Vol.22, No.7, pp.3084–3098, 2021.

180. Chameleon-inspired multifunctional plasmonic nanoplatforms for biosensing applications Yasamin Ziai1, Francesca Petronella, Chiara Rinoldi1, Pawel Nakielski, Anna Zakrzewska, Tomasz A. Kowalewski, Weronika Augustyniak, Xiaoran Li, Antonella Calogero, Izabela Sabala, Bin Ding, Luciano De Sio and Filippo Pierini, NPG Asia Materials vol. 14 no.18, 2022.

Prof. dr hab. T. A. Kowalewski Warszawa, March 4, 2022